OpenStudy (anonymous):

(x^2 + 2x - 1)(x^2 + 2x + 5)

OpenStudy (anonymous):

unfortunatly i dont think that is right lol, but thank you for helping

OpenStudy (anonymous):

you have to do 3 times 3 = 9 multiplications, so be prepared to write

OpenStudy (anonymous):

first multiply everything in the second parentheses by \(x^2\) then by \(2x\) and then by -1 finally combine like terms

OpenStudy (anonymous):

I just need a solution

hero (hero):

(x^2 + 2x - 1)(x^2 + 2x + 5) =x^2(x^2-2x-1) + 2x(x^2+2x-1)+5(x^2+2x-1) =x^4 -2x^3-x^2+2x^3+4x^2-2x+5x^2+10x-5 =x^4-2x^3+2x^3+4x^2-x^2+5x^2+10x-2x-5 =x^4+8x^2+8x-5

OpenStudy (anonymous):

if you just need a solution, why post here? use this http://www.wolframalpha.com/input/?i=%28x^2+%2B+2x+-+1%29%28x^2+%2B+2x+%2B+5%29 look at "alternate form" and it will be error free

OpenStudy (anonymous):

or use it to check the answer as you can see hero has it correct

hero (hero):

I don't have it correct. I made a mistake

hero (hero):

(x^2 + 2x - 1)(x^2 + 2x + 5) =x^2(x^2+2x-1) + 2x(x^2+2x-1)+5(x^2+2x-1) =x^4 +2x^3-x^2+2x^3+4x^2-2x+5x^2+10x-5 =x^4+2x^3+2x^3+4x^2-x^2+5x^2+10x-2x-5 =x^4+4x^3+8x^2+8x-5

OpenStudy (anonymous):

thank you!